# Surface to Seafloor: A Generative AI Framework for Decoding the Ocean Interior State

### v2

@@ 3.4 Statistics Combining Various Depths
Meridional Heat $\text{Meridional Heat Flux}=\rho c_{p}R\cos(\varphi)\int_{-1800~{}m}^{0~{}m}\int_{0^{\circ}}^{60^{\circ}}V(\lambda,\varphi,z,t)\,T(\lambda,\varphi,z,t)\,\,\,{\rm d}\lambda{\rm d}z.$ (9)
@@
[…]
@@
Figure 11: The instantaneous overturning stream function in depth coordinates. Here, we show the overturning stream function at each vertical level z and latitude for a fixed moment in time. The ocean simulation is compared to the AI mean and samples. The difference between the ocean simulation and the AI mean is shown in the top right. The pointwise standard deviation of the 100 generated samples is in the bottom right.
We see that the overall structure in the upper 400 meters of the domain is well captured, with errors increasing with depth. Furthermore, we see a larger discrepancy at lower latitudes and mid-depth. The largest standard deviation of the AI ensemble is at low latitudes and at depth. The lack of statistical equilibrium contributes to the error here. The largest error comes from the pool of cold water slowly finding its way along the seafloor, from northern latitudes towards the lower latitudes. This process takes time, and was not shown to the neural network during training. The samples of the generative AI method appear noisier than physically plausible, suggesting that a 3D UNet would be a better architecture for the present case. Regardless, the resulting variability is overestimated.
@@ 4 Conclusions
    While the model performs well in regions where surface observations strongly constrain subsurface dynamics, its predictive skill diminishes at greater depths and in areas where the ocean state is out-of-sample (because the ocean simulation is drifting over time) from ocean states similar to what has been seen previously.
@@
Future work will extend this approach to more complex ocean simulations that include seasonal cycles, bathymetry, and finer-scale turbulence, providing additional constraints on the generative model. Furthermore, integrating observational datasets with multiple modalities such as sea surface salinity, Argo float profiles, sea surface temperature, and velocity measurements from moorings will enable direct evaluation of the model’s performance in real-world ocean conditions, (Martin et al., 2025). Measurement uncertainty must also be accounted for. One avenue for doing so would be to draw from a distribution of consistent free-surface heights, effectively sampling from $\mathcal{P}(\text{subsurface}|\text{surface})\mathcal{P}_{s}(\text{surface})$ where $\mathcal{P}_{s}(\text{surface})$ incorporates uncertainty in the surface field. Incorporating theoretical predictions as additional “input” for the generative model will also be explored. This way, we can use a theory as a “physics-informed” guess, further refined through the generative method and a numerical ocean simulation with higher fidelity physics.

### v3

@@ 3.4 Statistics Combining Various Depths
$\text{Meridional Heat Flux}=\rho c_{p}R\cos(\varphi)\int_{-1800~m}^{0~m}\int_{0^{\circ}}^{60^{\circ}}V(\lambda,\varphi,z,t)\,T(\lambda,\varphi,z,t)\,\,\,{\rm d}\lambda{\rm d}z.$ (10)
@@
$\psi(\varphi,z,t)\stackrel{{\scriptstyle\textrm{def}}}{{=}}R\cos(\varphi)\int_{z}^{0}\int_{0^{\circ}}^{60^{\circ}}V(\lambda,\varphi,z^{\prime},t)d\lambda dz^{\prime}$ (11)
@@
Figure 11: The instantaneous overturning stream function in depth coordinates. Here, we show the overturning stream function at each vertical level z and latitude for a fixed moment in time. The ocean simulation is compared to the AI mean. The difference between the ocean simulation and the AI mean is shown in the third panel and the AI standard deviation in the last panel.
We see that the overall structure in the upper 400 meters of the domain is well captured, with errors increasing with depth. Furthermore, we see a larger discrepancy at lower latitudes and mid-depth. The largest standard deviation of the AI ensemble is at low latitudes and at depth. The lack of statistical equilibrium contributes to the error here. The largest error comes from the pool of cold water slowly finding its way along the seafloor, from northern latitudes towards the lower latitudes. This process takes time, and was not shown to the neural network during training. Regardless, the resulting variability is overestimated.
@@ 4 Conclusions
    While the model performs well in regions where surface observations strongly constrain subsurface dynamics, its predictive skill diminishes at greater depths and in areas where the ocean state is out-of-sample from ocean states similar to what has been seen previously.
@@
Future work will extend this approach to more complex ocean simulations that include seasonal cycles, bathymetry, and finer-scale turbulence, providing additional constraints on the generative model. Furthermore, integrating observational datasets with multiple modalities such as sea surface salinity, Argo float profiles, sea surface temperature, and velocity measurements from moorings will enable direct evaluation of the model’s performance in real-world ocean conditions, (Martin et al., 2025). Sparse measurements can also be incorporated as a part of the generative framework Babu et al. (2025). To properly incorporate satellite altimetry would be a space-time version of the “in-painting” exercise of that work, leading to a 4D probabilistic reconstruction.
Measurement uncertainty must also be accounted for. One avenue for doing so would be to draw from a distribution of consistent free-surface heights, effectively sampling from $\mathcal{P}(\text{subsurface}|\text{surface})\mathcal{P}_{M}(\text{surface})$ where $\mathcal{P}_{M}(\text{surface})$ incorporates uncertainty in the surface field. Incorporating theoretical predictions as additional “input” for the generative model will also be explored. This way, we can use a theory as a “physics-informed” guess, further refined through the generative method and a numerical ocean simulation with higher fidelity physics. For example, one could use theory to provide further conditional information to augment the prediction of the generative method. Symbolically we would be drawing from the distribution $\mathcal{P}(\text{subsurface fields}|\text{surface fields},\text{theory}),$ which could possibly further reduce ensemble spread in samples.
@@ Acknowledgements.
$\mathcal{C}[\mathcal{F}]=\int\mathcal{P}_{J}(\text{subsurface fields},\text{surface fields})\|\mathcal{F}(\text{surface fields})-\text{subsurface fields}\|^{2}$ (12)
[…]
$\mathcal{P}_{J}(x,y)=\frac{1}{N}\sum_{n=1}^{N}\delta(x-x^{n})\delta(y-y^{n})$ (13)
as our explicit data-distribution, where $\delta$ is the Dirac-delta function; however, a well-trained neural networks will regularize the above data-distribution so that it may generalize to unseen data-pairs. Ideally, the regularized distribution should be related to the “true distribution” of the underlying numerical simulation.
Our cost functional with the same notation is
$\mathcal{C}[\mathcal{F}]=\int\mathcal{P}_{J}(x,y)\|\mathcal{F}(x)-y\|^{2}.$ (14)
To minimize the above functional across all functions $\mathcal{F}$ we take the variational derivative of the above expression with a small function perturbation $\delta\mathscr{F}(x)$ and neglect higher order terms to yield
$\delta\mathcal{C}=\mathcal{C}[\mathcal{F}+\delta\mathscr{F}]-\mathcal{C}[\mathcal{F}]\approx 2\int\mathcal{P}_{J}(x,y)\delta\mathscr{F}(x)\cdot\left(\mathcal{F}(x)-y\right)$ (15)
We now use the identity $\mathcal{P}_{J}(x,y)=\mathcal{P}_{M}(x)\mathcal{P}(y|x)$ to rewrite the above expression as
$\int\mathcal{P}_{J}(x,y)\delta\mathscr{F}(x)\cdot\left(\mathcal{F}(x)-y\right)=\int dx\mathcal{P}_{M}(x)\delta\mathscr{F}(x)\cdot\int dy\mathcal{P}(y|x)\left(\mathcal{F}(x)-y\right).$ (16)
Carrying out the integral over $y$ we identify
$\int dy\mathcal{P}(y|x)\mathcal{F}(x)=\mathcal{F}(x)\text{ and }\int dy\mathcal{P}(y|x)y=\mathbb{E}_{y}\mathcal{P}(y|x),$ (17)
which reduces our functional derivative to
$\delta\mathcal{C}=2\int dx\mathcal{P}_{M}(x)\delta\mathscr{F}(x)\cdot\left(\mathcal{F}(x)-\mathbb{E}_{y}\mathcal{P}(y|x)\right).$ (18)
Setting $\delta\mathcal{C}=0$ and noting that this must hold across all variations $\delta\mathscr{F}$ yields the expression
$\mathcal{F}(x)=\mathbb{E}_{y}\mathcal{P}(y|x),$ (19)
i.e., the minimizer is the conditional expectation.
At this point it is worth contrasting the above expression with what is obtained by a generative method. The generative model directly tries to capture (and sample from) the conditional distribution $\mathcal{P}(y|x)$ whereas the deterministic method calculates directly the expected value $\mathcal{F}(x)=\mathbb{E}_{y}\mathcal{P}(y|x)$. Thus to compare generative models to traditional models (minimizing with respect to MSE) we generate samples from the distribution $\mathcal{P}(y|x)$ and take an empirical average afterwards.
There are, however, practical issues that may prevent such an identification of traditional machine learning methods with generative methods. One issue comes from different architectures or forms of implicit regularization that are used when training neural networks. Another issue is that the one may require a large number of empirical samples to generate a particular statistic. If either model overfits on the training dataset, then these results do not apply since we the connection between the two only happens for the “global” minimizer of the cost functional for the underlying data-distribution (as opposed to the empirical data-distribution).